OpenStudy (anonymous):

What is the most precise name for quadrilateral ABCD with vertices A(-4,-1), B(-2,-5), C(4,-2) and D(2,2)? Parallelogram rhombus quadrilateral

OpenStudy (anonymous):

So, I used the midpoint formula. M = x1 + x2/2, y1 + y2/2 AB M = (-1 + -2/2, -1 + -5/2) = (-1.5,-3) BC M = (-2 + 4/2, -5 + -2/2) = (1,-3.5) CD M = (4+ 2/2, -5 + -2/2) = (3,0) DA M = (2 + -4/2, 2 + -1/2) = (-1,0.5)

OpenStudy (anonymous):

I forgot the fourth option, rectangle

OpenStudy (mathmath333):

is the forth option \(\huge kite\)

OpenStudy (anonymous):

no

OpenStudy (mathmath333):

rectangle ?

OpenStudy (anonymous):

there is one more step to the problem to find out but I can't figure it out

OpenStudy (anonymous):

one sec I will send a picture of that page in the book

OpenStudy (anonymous):

Does that help?

OpenStudy (mathmath333):

wait

OpenStudy (anonymous):

How do you tag people on here?

OpenStudy (mathmath333):

@School_is_hard

OpenStudy (anonymous):

oh ok thank you

OpenStudy (mathmath333):

well at the previous it is written in the question to find the midpoint , but in your posted question ,it is not so, so no need to find midpoint here

OpenStudy (anonymous):

oh ok, so how do I find out which answer it is?

OpenStudy (mathmath333):



OpenStudy (anonymous):

I know what they look like. But I don't think they are wanting me to go by looks. because in the book a parallelogram would be a slope of -1, A triangle classifying isosceles, scalene, equilateral, and thing like that

OpenStudy (mathmath333):

quadrilateral can be any figure with four sides 

OpenStudy (mathmath333):

so what u think what is the answer

OpenStudy (mathmath333):

do u have to right a paragraph or something as long for answer

OpenStudy (anonymous):

ether a quadrilateral or a rombus

OpenStudy (anonymous):

i meant rhombus

OpenStudy (anonymous):

no its multiple choice

OpenStudy (mathmath333):

\(\huge no\)

OpenStudy (anonymous):

and I have two like this one

OpenStudy (mathmath333):

do u think in the figure the opposite sides are equal

OpenStudy (anonymous):

no

OpenStudy (mathmath333):

it is the red figure right ?

OpenStudy (anonymous):

no the green

OpenStudy (anonymous):

look at the page I sent page 402 explains it

OpenStudy (mathmath333):

in the question the co-ordinates u gave only matches with the red figure

OpenStudy (anonymous):

yes but the green is what I got from finding the midpoints

OpenStudy (anonymous):

And on page 402 you have to find the midpoint then distance of the green to classify the figure

OpenStudy (mathmath333):

but i said earlier that u didnot need to find the midpoint here like the previous question as it is not written in the question to find the midpoint